OpenStudy (anonymous):

Please help Solve for x sin(2x)=cos(3x) I did it by expanding everything with double angle and addition identities but I ended up with 2sin(x)+2sin^2(x)-cos(x)=-1 Not sure where to go or if there is an easier way to do it

OpenStudy (danjs):

hmm, lots of properties to use prolly

OpenStudy (danjs):

sin(2x)=cos(3x) 2*sin(x)*cos(x) = cos(2x + x) , what was the cox(x+y)= thing?

OpenStudy (danjs):

i think i have it, if you return

OpenStudy (anonymous):

Sorry I left But anyway that's what I did and I ended up with 2sin(x)+2sin^2(x)-cos(x)=-1

OpenStudy (anonymous):

And I did that by writing out the result of the cos(2x+x) and getting cos^2(x)-cos(x)-2sin^2(x)cos(x) I then factored out the cos(x) since it was a common factor On the LHS I got 2sin(x)cos(x) and then I divide both sides by cos(x) Getting 2sin(x)=cos(x)-1-2sin^2(x) Then I left the -1 on the RHS and put the other two terms onto the LHS Getting 2sin(x)+2sin^2(x)-cos(x)=-1

OpenStudy (danjs):

2*sin(x)*cos(x) = cos(2x + x) 2*sin(x)*cos(x) = cos(2x)*cos(x) - sin(2x)*sin(x) 2*sin(x)*cos(x) = ( 1 - 2sin^2(x))*cos(x) - 2*sin(x)*cos(x)*sin(x) 2*sin(x)*cos(x) = cos(x) - 2sin^2(x)*cos(x) - 2sin^2(x)*cos(x) 2sin(x) = 1 - 4sin^2(x) i think that is it,, solve quadratic

OpenStudy (anonymous):

Yeah thanks I'll try it

OpenStudy (danjs):

not sure if it supposed to be a 'nice' angles or not

OpenStudy (anonymous):

No it's not They're supposed to be complicated lol

OpenStudy (danjs):

4u^2 + 2u - 1 = 0

OpenStudy (danjs):

sin(x) = ......

OpenStudy (danjs):

yeah no,, goodluck

OpenStudy (anonymous):

Um just checking that's supposed to be 3 and not 4 right?

OpenStudy (danjs):

i was just looking over the linked sheet, the cofunctions formula... sin(x)=cos(pi/2 - x) maybe replace the sin(2x) with cos(pi/2 - 2x) cos(pi/2 - 2x) = cos(3x)

OpenStudy (danjs):

pi/2 - 2x = 3x

OpenStudy (danjs):

if that is legal, .. and cos(angle) is symmetric for + or - angle, so set that to +3x and -3x

OpenStudy (danjs):

i think i did this prob before on here sometime, or something like it

OpenStudy (danjs):

@brill no, 4, 3rd line in the thing, you distribute, and have -2 and a -2

OpenStudy (anonymous):

Oh yeah you're right, forgot the two whoops

OpenStudy (anonymous):

Thanks! I solved it using quadratic formula.

OpenStudy (danjs):

cool, check to see if that second method also is the same answer... brb in a few

OpenStudy (anonymous):

Yeah I tried it but the most I could take it is a -sin(x)=cos(x)-4sin^2(x)cos(x)